# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

### v1

@@ Abstract
Index Terms:
Communication-centric ISAC, opportunistic sensing, OTFS modulation, LEO satellites, space targets, high-speed targets, range migration, maximum likelihood estimation.
@@ I Introduction
Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [6, 7]. OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing better communication performance in doubly-dispersive fading channels as compared to schemes using orthogonal frequency division multiplexing [8, 9, 10]. They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate [^1] In this study, range-rate refers to the radial component of velocity, i.e., the rate at which the distance between the object and the observer is changing. of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12]. These unique features enable OTFS-based LEO satellites to inherently support ISAC in various forms [13, 14]. Unlike terrestrial stations, LEO-based stations have the capability to also cover aerial and space regions, and this expanded coverage opens up new application scenarios, such as real-time environmental monitoring, enhanced global navigation systems, and space exploration. Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18]; this approach is indeed considered in this study.
@@
Previous studies assume that the length (range) of a channel path between the transmitter and receiver remains constant within the OTFS frame, which is typically valid for terrestrial applications. However, this assumption may not hold true for satellite-based space sensing, as the significant relative motion between the transmitter, monitored target, and receiver can cause range migration during the OTFS frame. With a transmitter on a LEO satellite and a radar receiver on the same or a different satellite, or an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecraft, and unresponsive satellites) may have a range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a range-rate of $\bar{v}=15$ km/s, the range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the range resolution $c/(2M\Delta)\approx 19.5$ m, where $c$ is the speed of light.
@@
  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive delay-Doppler search. First, a block orthogonal matching pursuit (BOMP) algorithm [35] is used to obtain coarse information on the initial range and range-rate. Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower delay-Doppler region. The proposed procedure can be combined with the CLEAN algorithm in [36, 37, 38, 39, 15] to handle multiple targets.
@@
In the following, $\mathbb{R}$ and $\mathbb{C}$ are the set of real and complex numbers, respectively. Column vectors and matrices are denoted by lowercase and uppercase boldface letters, respectively. The symbols $(\,\cdot\,)^{*}$, $(\,\cdot\,)^{\mathsf{T}}$, and $(\,\cdot\,)^{\mathsf{H}}$ denote conjugate, transpose, and conjugate-transpose, respectively. $\bm{I}_{M}$ is the $M\times M$ identity matrix. $(\bm{a})_{i}$ denotes the $i$-th entry of the vector $\bm{a}$. $(\bm{A})_{i,j}$ denotes the entry in the $i$-th row and $j$-th column of the matrix $\bm{A}$. $(\bm{A}_{1}\ \cdots\ \bm{A}_{n})$ and $(\bm{A}_{1};\ \cdots\ ;\bm{A}_{n})$ denote the horizontal and vertical concatenations of the matrices $\bm{A}_{1},\ldots,\bm{A}_{n}$. $\mathbbm{1}_{\mathcal{A}}$ is the indicator function of the condition $\mathcal{A}$, i.e., $\mathbbm{1}_{\mathcal{A}}=1$, if $\mathcal{A}$ holds true, and $\mathbbm{1}_{\mathcal{A}}=0$ otherwise. $[\,\cdot\,]_{M}$ denotes the mod $M$ operation. $\delta(t)$ and $\delta[n]$ denote the Dirac delta function and the discrete unit sample function, respectively. ${}\Re\left\{\cdot\right\}$ denotes the real part. Finally, $\mathrm{i}$, $\odot$, and $\mathrm{E}[\,\cdot\,]$ denote the imaginary unit, the Schur product, and the statistical expectation, respectively.
@@ II System description
Consider a sensing system consisting of a communication transmitter and a radar receiver. The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates with a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and cannot be modified to adjust the radar needs. The radar receiver aims to estimate the initial range (delay), range-rate (Doppler shift), and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellites, an aircraft, a high altitude platform, or a terrestrial station. The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15], whereas, for a bistatic configuration, this requires the presence of some coordination [^4] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link. or a separate reference channel able to demodulate the communication signal [40, 41]. For a more illustration, we consider a bistatic scenario, as shown in Fig. 1, but the following developments similarly apply to the monostatic case. We initially assume the presence of a single target to simplify exposition.
@@ II-A Transmit signal
[…]
@@
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$. By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$. Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $[-\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $[-\Delta/2,M\Delta-\Delta/2]$. [^6] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $[-\Delta/2,M\Delta-\Delta/2]$ is present. In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$.
@@ II-B Design assumptions
During an OTFS frame, the target, the communication transmitter, and the radar receiver can be assumed to move with a constant velocity. The position of the target at time $t\in[0,NT]$ is $\bm{p}(t)=\bm{p}-\bm{v}t\in\mathbb{R}^{3}$, where $\bm{p}$ and $\bm{v}$ are the initial position and the velocity, respectively; similarly, the positions of communication transmitter and the radar receiver at time $t\in[0,NT]$ are $\bm{p}_{\rm tx}(t)=\bm{p}_{\rm tx}-\bm{v}_{\rm tx}t\in\mathbb{R}^{3}$ and $\bm{p}_{\rm rx}(t)=\bm{p}_{\rm rx}-\bm{v}_{\rm rx}t\in\mathbb{R}^{3}$, respectively. Accordingly, for $t\in[0,NT]$, the (bistatic) target range is
@@ II-C Received signal
is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$). The support of $y(t)$ is contained in the time interval $[0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $[-\Delta/2,M\Delta+v_{\max}-\Delta/2]$.
@@
[…]
@@ II-D OTFS demodulation
[…]
@@
×$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
@@
×$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
×$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
@@
×$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
@@ III Ideal shaping filters
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (23)
@@
(($\star$)( ⋆ ) $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
×$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
@@
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
@@
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (25)
@@ III-A Some insights on Proposition 2
Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 00 of width $2/Q$ (between zero crossings) and $Q-2$ sidelobes of width $1/Q$ (between zero crossings) in $[-0.5,0.5]$. The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [42]; following [8], for large $Q$, we can assume $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ is a design parameter that depends upon the desired approximation error. Accordingly, the following remarks are in order.
@@
For large $N/B$ and $M$, the sums in (26) only contain few dominant terms. To illustrate this point, write first the range at time $bT_{B}$ and the range-rate as
@@
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (28)
Finally, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Phi[k,l]|\approx 0$ if $[k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $[l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4. ∎

### v2

@@ I Introduction
Orthogonal time frequency space (OTFS) modulation is a promising transmission scheme for future LEO constellations [6, 7]. OTFS-based transceivers exploit channel diversity across both time and frequency, thus providing better communication performance in doubly-dispersive fading channels as compared to schemes using orthogonal frequency division multiplexing [8, 9, 10]. They are also suitable for radar sensing, as the underlying signal processing resembles that of pulsed-Doppler radars [11]; in particular, the range and range-rate of prospective scattering objects can be inferred from the estimated delay-Doppler channel response [12]. These unique features enable OTFS-based LEO satellites to inherently support ISAC in various forms [13, 14]. Unlike terrestrial stations, LEO-based stations have the capability to also cover aerial and space regions, and this expanded coverage opens up new application scenarios, such as real-time environmental monitoring, enhanced global navigation systems, and space exploration. Communication-centric ISAC solutions appear particularly promising in this context, as existing data signals can be opportunistically reused to add the sensing functions at a low cost, without altering the underlying communication protocol or affecting its performance [12, 15, 16, 17, 18, 19]; this approach is indeed considered in this study.
@@
Previous studies assume that the length (range) of a channel path between the transmitter and receiver remains constant within the OTFS frame, which is typically valid for terrestrial applications. However, this assumption may not hold true for satellite-based space sensing, as the significant relative motion between the transmitter, monitored target, and receiver can cause range migration during the OTFS frame. With a transmitter on a LEO satellite and a radar receiver on an aircraft at a lower altitude, non-cooperative space targets (e.g., space debris, asteroids, spacecrafts, missiles, and unresponsive satellites) may have a bistatic range-rate of several km/s; for a system with $M=512$ subcarriers, $N=128$ symbol intervals, a subcarrier spacing of $\Delta=15$ kHz, a symbol interval of $T=1/\Delta$, and a bistatic range-rate of $\bar{v}=15$ km/s, the bistatic range variation in the OTFS frame is $\bar{v}NT=128$ m and greatly exceeds the bistatic range resolution $c/(M\Delta)\approx 39$ m, where $c$ is the speed of light.
@@
  • We propose an approximate implementation of the maximum likelihood (ML) estimator for target parameters, which avoids exhaustive two-dimensional search in the delay-Doppler plane. First, a block orthogonal matching pursuit (BOMP) algorithm [36] is used to obtain coarse information on the initial range and range-rate. Next, a refined estimate of the target parameters is derived using a bank of matched filters focused on a narrower delay-Doppler region. The proposed procedure can be combined with the CLEAN algorithm in [37, 38, 39, 40, 16] to handle multiple targets.
@@
In the following, $\mathbb{R}$ and $\mathbb{C}$ are the set of real and complex numbers, respectively. Column vectors and matrices are denoted by lowercase and uppercase boldface letters, respectively. The symbols $(\,\cdot\,)^{*}$, $(\,\cdot\,)^{\mathsf{T}}$, and $(\,\cdot\,)^{\mathsf{H}}$ denote conjugate, transpose, and conjugate-transpose, respectively. $\bm{I}_{M}$ is the $M\times M$ identity matrix. $(\bm{a})_{i}$ denotes the $i$-th entry of the vector $\bm{a}$. $(\bm{A})_{i,j}$ denotes the entry in the $i$-th row and $j$-th column of the matrix $\bm{A}$. $(\bm{A}_{1}\ \cdots\ \bm{A}_{n})$ and $(\bm{A}_{1};\ \cdots\ ;\bm{A}_{n})$ denote the horizontal and vertical concatenations of the matrices $\bm{A}_{1},\ldots,\bm{A}_{n}$. $\mathbbm{1}_{\mathcal{A}}$ is the indicator function of the condition $\mathcal{A}$, i.e., $\mathbbm{1}_{\mathcal{A}}=1$, if $\mathcal{A}$ holds true, and $\mathbbm{1}_{\mathcal{A}}=0$ otherwise. $\cdot\,]_{M}$ denotes the mod $M$ operation. $\delta(t)$ and $\delta[n]$ denote the Dirac delta function and the discrete unit sample function, respectively. $\Re\left\{\cdot\right\}$ denotes the real part. Finally, $\mathrm{i}$, $\odot$, and $\mathrm{E}[\,\cdot\,]$ denote the imaginary unit, the Schur product, and the statistical expectation, respectively.
@@ II System description
Consider a sensing system consisting of a communication transmitter and a radar receiver. The communication transmitter is mounted on a LEO satellite, uses an OTFS modulation, and operates with a carrier frequency $c/\lambda$, where $\lambda$ is the carrier wavelength; the emitted waveform is specified by the underlying communication protocol and cannot be modified to adjust the radar needs. The radar receiver aims to estimate the initial range (delay), range-rate (Doppler shift), and amplitude of high-speed targets illuminated by the communication transmitter; it may be mounted on the same platform hosting the communication transmitter (monostatic configuration) or on a separate device (bistatic configuration), e.g., another LEO satellites, an aircraft, a high altitude platform, or a terrestrial station. The receiver is synchronized with the transmitter and knows the emitted symbols: for a monostatic configuration, this is readily verified [15, 16], whereas, for a bistatic configuration, this requires the presence of some coordination [^3] For example, when the radar receiver is on a different LEO satellite, the coordination may be established via an inter-satellite link. or a separate reference channel able to demodulate the communication signal [41, 42, 43].
For illustration, we consider a bistatic radar configuration, as depicted in Fig. 1. In this setup, the bistatic range of a target is defined as the total distance traveled by the probing signal (first from the communication transmitter to the target, and then from the target to the radar receiver). Correspondingly, the bistatic range-rate refers to the time derivative of the bistatic range, i.e., the rate at which the bistatic range changes due to motion of the transmitter, the target, or the receiver. It is important to note that the following analysis also applies to the monostatic case, where the transmitter and receiver are co-located; in such a configuration, the bistatic range and range-rate simplify to twice the monostatic range and range-rate, respectively. To streamline the exposition, we will omit the adjective "bistatic" from this point forward and use the terms range and range-rate to refer to the bistatic quantities unless otherwise specified. Finally, we begin by assuming the presence of a single target to simplify the initial discussion.
@@ II-A Transmit signal
Consider an OTFS frame with bandwidth $M\Delta$ and duration $NT$, where $M$ and $N$ are the number of subcarriers and symbol intervals, respectively, $\Delta$ is the subcarrier spacing, and $T=1/\Delta$ is the symbol interval, while $1/(M\Delta)$ and $1/(NT)$ are the delay and Doppler resolutions, respectively. Up to a positive scaling factor accounting for the radiated energy, the baseband signal emitted by the communication transmitter is
@@
for $n=0,\ldots,N-1$ and $m=0,\ldots,M-1$. By considering $g_{\rm tx}(t)$ time limited in $[0,T]$, then $x(t)$ has support $[0,NT]$. Also, by assuming the spectrum of $g_{\rm tx}(t)$ contained in $\Delta/2,\Delta/2]$, then the spectrum of $x(t)$ is contained in $\Delta/2,M\Delta-\Delta/2]$. [^4] Since a time limited signal cannot be strictly band limited, some spectral leakage outside $\Delta/2,M\Delta-\Delta/2]$ is present. In practice, the transmit shaping filter may not be exactly zero outside $[0,T]$ to reduce such spectral leakage; in this latter case, $x(t)$ is only approximately contained in $[0,NT]$.
@@ II-B Design assumptions
During an OTFS frame, the target, the communication transmitter, and the radar receiver can be assumed to move with a constant velocity. The position of the target at time $t\in[0,NT]$ is $\bm{p}(t)=\bm{p}-\bm{v}t\in\mathbb{R}^{3}$, where $\bm{p}$ and $\bm{v}$ are the initial position and the velocity, respectively; similarly, the positions of communication transmitter and the radar receiver at time $t\in[0,NT]$ are $\bm{p}_{\rm tx}(t)=\bm{p}_{\rm tx}-\bm{v}_{\rm tx}t\in\mathbb{R}^{3}$ and $\bm{p}_{\rm rx}(t)=\bm{p}_{\rm rx}-\bm{v}_{\rm rx}t\in\mathbb{R}^{3}$, respectively. Accordingly, for $t\in[0,NT]$, the target range is
@@ II-C Received signal
is the baseband representation of the target echo with respect to the reference frequency $\frac{c+\bar{v}_{\min}}{\lambda}$ and $r(t)=\bar{r}(t)-\bar{r}_{\min}$ is the excess target range (i.e., the target range in excess with respect to $\bar{r}_{\min}$). The support of $y(t)$ is contained in the time interval $0,NT+r_{\max}/c]$, and its spectrum is negligible outside the frequency interval $\Delta/2,M\Delta+v_{\max}/\lambda-\Delta/2]$.
@@
$=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}h_{n}(\tau,\nu)x_{n}\left(t-\tau\right)\mathrm{e}^{\mathrm{i}2\pi\nu(t-\tau)}d\tau d\nu$ (12)
@@ II-D OTFS demodulation
$\gamma(\tau,\nu)=\int_{-\infty}^{\infty}g_{\rm tx}(\beta)g^{*}_{\rm rx}(\beta-\tau)\mathrm{e}^{-\mathrm{i}2\pi\nu(\beta-\tau)}d\beta$ (15)
@@
$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)g^{*}_{\rm rx}\!\left(t-nT\right)$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi m^{\prime}\Delta\left(t-n^{\prime}T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{-\mathrm{i}2\pi m\Delta\left(t-nT\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}t}dt$
@@
$\quad\times\alpha\mathrm{e}^{\mathrm{i}2\pi m\Delta\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\left(n^{\prime}T+\frac{r\left(n^{\prime}T\right)}{c}\right)}$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c}\right)}$
$\quad\times\int_{-\infty}^{\infty}g_{\rm tx}\!\left(\beta\right)g^{*}_{\rm rx}\!\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left((m-m^{\prime})\Delta-\frac{v}{\lambda}\right)\left(\beta-(n-n^{\prime})T+\frac{r\left(n^{\prime}T\right)}{c}\right)}d\beta$
@@
$\quad\times\gamma\left((n-n^{\prime})T-\frac{r\left(n^{\prime}T\right)}{c},(m-m^{\prime})\Delta-\frac{v}{\lambda}\right),$
@@ III Ideal shaping filters
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (23)
@@
$(\star)$ $=M\alpha\sum_{n=0}^{N-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)n}$
$\quad\times\left[\frac{1}{M}\sum_{m=0}^{M-1}\mathrm{e}^{\mathrm{i}2\pi\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)m}\right]$
@@
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(nT\right)\Delta}{c}\right)\right)$
@@
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right)$ (25)
@@ III-A Some insights on Proposition 2
Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5]. The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [44]; following [8], for large $Q$, we can assume $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ is a design parameter that depends upon the desired approximation error. Accordingly, the following remarks are in order.
@@
For large $N/B$ and $M$, the sums in (26) only contain few dominant terms. To illustrate this point, denote by $R_{\rm r}=c/(M\Delta)$ and $R_{\rm rr}=\lambda/(NT)$ the range and range-rate resolutions [41], respectively, and write the range at time $bT_{B}$ and the range-rate as
@@
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l_{b,{\rm int}}-l_{b,{\rm fra}}}{M}\right)\right).$ (28)
Finally, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4. ∎